OpenStudy (anonymous):

Find the volume of a cylinder with a height of 17 meters and a circular base with a diameter of 16 meters.

OpenStudy (angela210793):

V=B*l=pi*8^2*17

OpenStudy (angela210793):

B=area of base=pi R^2

OpenStudy (anonymous):

answer is 68 or no

OpenStudy (anonymous):

no.....

OpenStudy (angela210793):

oops...

OpenStudy (anonymous):

volume = π*r^2*h = 3.14*(8)^2*17

OpenStudy (angela210793):

u answer Harkirat......my brain's on vacation.....:

OpenStudy (anonymous):

213.52??

OpenStudy (anonymous):

it is 3416.327

OpenStudy (anonymous):

does ^ mean divide??

OpenStudy (anonymous):

^ means "raised to the power" so ^2 means square i.e. 5^2 = 5x5 =25 volume = π*r^2*h = 3.14*(8)^2*17 = 3.14 * 64 * 17 = 3416.32

OpenStudy (anonymous):

hope it is clear now ..........

OpenStudy (anonymous):

oh ok I didnt now what that meant thnx

OpenStudy (anonymous):

from now on keep in mind that ^ means raised to the power and the number following it gives the power ^2 means number squared ^3 means number cubed

OpenStudy (anonymous):

got a medal for me ... o( :})

OpenStudy (angela210793):

U got 2 medals and I didn't get even 1? :S :P

OpenStudy (anonymous):

here's one for you {{ :))

OpenStudy (angela210793):

lol...thnx :)